Write an equation in point-slope form of the line that passes through
the points (-3,-4) and (3,4).

Answer by mananth(16946)

Write an equation in point-slope form of the line that passes through
the points (-3,-4) and (3,4).
..
slope m= y2-y1/x2-x1
m=4-(-4)/3-(-3)
m=8/6 = 4/3
...
Take point (-3,-4) & m=4/3
find b in equation y= mx+b
-4 = -3*4/3 +b
-4=-4+b
b=0
..
The equation:
y=4x/3
...
m.ananth@hotmail.ca